1) The vapor pressure of liquid cobalt is 400. mm Hg at 3.03×103 K. Assuming that ΔHvap for Co (450 kJ/mol) does not change significantly with temperature, calculate the vapor pressure of liquid Co at 3.07×103 K.

2) The vapor pressure of methyl acetate (CH3COOCH3) has been measured over a range of temperatures.

(a) Which of the following should give a linear plot (temperatures in K)?

On the x-axis: _____T or lnT or 1/T   On the y-axis: _____P or lnP 1/P



(b) When the appropriate plot is made, the resulting straight line has a slope of -4.20×103 K.

Use this information to determine ΔHvap for CH3COOCH3. ______kJ/mol
Answer #1

Q1. vapor pressure of liquid Co at 3.7 x 103 K = 504.8 mmHg

Q2. Hvap for CH3COOCH3 = 34.9 kJ/mol

Explanation

Q1. According to Clausius - Clapeyron equation,

ln(P2 / P1) = (Hvap / R) * (1/T1 - 1/T2)

ln(P2 / 400 mmHg) = (450 x 103 J/mol / 8.314 J/mol-K) * (1/3030 K - 1/3070 K)

ln(P2 / 400 mmHg) = 0.233

P2 / 400 mmHg = e0.233

P2 / 400 mmHg = 1.262

P2 = (400 mmHg) * (1.262)

P2 = 504.8 mmHg